# Climate-economy projections under shared socioeconomic pathways and net-zero scenarios

### v1

@@ Abstract
We examine future trajectories of the social cost of carbon, global temperatures, and carbon concentrations using the cost-benefit Dynamic Integrated Climate-Economy (DICE) model calibrated to the five Shared Socioeconomic Pathways (SSPs) under two mitigation scenarios: achieving net-zero carbon emissions by 2050 and by 2100. The DICE model is calibrated to align industrial and land-use carbon emissions with projections from six leading process-based integrated assessment models (IAMs): IMAGE, MESSAGE–GLOBIOM, AIM/CGE, GCAM, REMIND–MAgPIE and WITCH–GLOBIOM. We find that even with aggressive mitigation (net-zero by 2050), global temperatures are projected to exceed $2^{\circ}\text{C}$ above preindustrial levels by 2100, with estimates ranging from $2.5^{\circ}\text{C}$ to $2.7^{\circ}\text{C}$ across all SSPs and IAMs considered. Under the more lenient mitigation scenario (net-zero by 2100), global temperatures are projected to rise to between $3^{\circ}\text{C}$ and $3.7^{\circ}\text{C}$ by 2100. Additionally, the social cost of carbon is estimated to increase from approximately USD 30–50 in 2025 to USD 250–400 in 2100.
Keywords: Climate change, social cost of carbon, integrated climate-economy modelling, shared socioeconomics pathways, representative concentration pathways, net-zero emission scenarios
@@ 1 Introduction
The impacts of climate change on global warming have become increasingly more evident worldwide, prompting the negotiation of the Paris Agreement in 2015 by 196 parties United Nations Treaty Collection (2015). This landmark treaty aims to limit the global temperature increase to $2^{\circ}\text{C}$ above preindustrial levels, with a preferred target of $1.5^{\circ}\text{C}$. To achieve these goals, many countries, cities, businesses, and other institutions have committed to achieving net-zero carbon emissions—where the greenhouse gases released into the atmosphere are balanced by those removed. Hereafter, “carbon emission” refers to the emission of all greenhouse gases. Over 140 countries, including China, India, and the European Union countries, have set targets to reach net-zero emissions by around 2050. This has resulted in more public and private funds being provided for green projects involving renewable energy and energy conservation, as the world works to mitigate the effects of global climate change.
@@
Since the early 1990s, Integrated Assessment Models (IAMs) have become a central analytical tool in climate change economics. See reviews in Weyant (2017). The mitigation and climate impact assessments in the Intergovernmental Panel on Climate Change (IPCC) Assessment Reports [^1] The IPCC is the United Nations body responsible for assessing the science related to climate change. Its latest report, the Sixth Assessment Report (2023), is available at www.ipcc.ch. have primarily relied on IAMs developed by various research groups. These models can be broadly categorized into two types: cost-benefit IAMs, which estimate optimal mitigation strategies by balancing economic costs and climate damages, and process-based IAMs, which form the basis of IPCC’s assessments of transformation pathways towards temperature targets. However, most cost-benefit IAMs have historically relied on a single reference pathway and have not systematically incorporated the full range of Shared Socioeconomic Pathway (SSP) scenarios. As a result, there is a lack of integrated research that quantitatively aligns the baseline assumptions of process-based IAMs with cost-benefit modelling frameworks under alternative net-zero emission trajectories.
@@
In this paper, we study the question whether zero emission targets can lead to a temperature increase under $2^{\circ}\text{C}$ by 2100 or more ambitious policies should be adopted. We study this using the cost-benefit dynamic integrated climate-economy (DICE) model calibrated to baseline SSPs and emissions from the leading process-based models. DICE was orginally proposed more than 30 years ago in Nordhaus et al. (1992) and was regularly refined with the latest beta version update in 2023222The most recent version of the DICE model is available at https://williamnordhaus.com%https://williamnordhaus.com/dicerice-models.. It is cost-benefit IAM that estimates optimal mitigation levels relative to the economic costs of climate impacts. DICE model is one of the three main IAMs used by the United States government to determine the social cost of carbon; see Interagency Working Group on Social Cost of Greenhouse Gases (2016). The other two are PAGE (Policy Analysis of the Greenhouse Effect model, see Hope (2008)) and FUND (Climate Framework for Uncertainty, Negotiation, and Distribution model, see Anthoff and Tol (2010)). Even though the DICE model has limitations in the model structure and model parameters debated in the literature Grubb et al. (2021); Pindyck (2017), it has become the typical reference point for climate-economy modelling, and is used in our study.
@@
Almost any process-based IAM can be run for different SSPs. Among these, for each SSP a single IAM interpretation was selected as the so-called representative marker scenario for recommended use by future analyses of climate change, its impacts, and response measures. These are referred as marker SSPs, discussed in detail in e.g. Riahi et al. (2017) and defined in Section 2.2. For example, marker SSP1 corresponds to SSP1 run using IMAGE. The selection of markers was guided by two main considerations: the internal consistency of the full set of SSP markers, and the ability of the different models to represent distinct characteristics of the storylines. The SSP data and process-based IAMs forecasts are available from SSPdata-IIASA. This database does not include cost-benefit IAMs such as DICE. Even though marker SSPs are standard references, non-marker SSPs are also considered to be important since they provide insights into possible alternative scenario interpretations of the same basic SSP elements and storylines. In our study, we calculate results for DICE model calibrated to baseline scenarios of all five marker SSPs; we also calculate non-marker scenarios when IAM data are available for all five SSPs.
@@
Our study is motivated by paper Yang et al. (2018) that examines DICE results under SSP1–SSP5, where the results of the China Climate Change integrated assessment model (C3IAM) were used to re-estimate some parameters in the DICE model. In our study, we re-calibrate the DICE model to match the SSP baseline scenarios [^4] The baseline SSP scenarios are the reference cases for mitigation, climate impacts and adaptation analyses. They represent future developments in the absence of new climate policies and mitigation scenarios. and IAMs forecasts for the emission of greenhouse gases. Then, re-calibrated DICE model is used to produce scenarios of future for climate and economy variables under SSP1–SSP5 and assuming scenarios for carbon emission reduction leading to net-zero carbon emission in 2050 and 2100 (we also calculated scenarios for zero industrial emission only). The main observation is that even if a zero emission target is achieved worldwide by 2050, the temperature increased by 2100 will be larger than 2C in the range 2.5-2.7C. For this scenario, the social cost of carbon is increasing about 10 times from USD 30-50 in 2025 to USD 250-400 in 2100 depending on SSP-IAM. This provides a carbon price benchmark for policy makers and is generally consistent with Riahi et al. (2022) and Yang et al. (2018).
We made two contributions. First, we systematically calibrated DICE to the full set of baseline SSP pathways, utilising outputs from the main six process-based integrated assessment models rather than a single reference scenario. This provides a more comprehensive outlook on socio-economic uncertainty and potential variations in emissions trajectories. Second, introducing net-zero emission (and zero industrial emission) scenarios within this calibrated framework provides policy-relevant insights on the potential to achieve stringent temperature targets, such as the 1.5–2°C target set out in the Paris Agreement.
@@ 2 Methodology
In this section, we outline the main DICE model equations and the calibration of the DICE exogenous functions to the SSP-IAM data. The results for climate-economy trajectories under DICE calibrated to different SSPs-IAMs will be presented in the next section.
@@ 2.1 Model
In our study, we calculate the climate-economy projections using DICE-2016. Specifically, we use the version DICE2016R-091916ap.gms. We note that beta 2023 version is available but we prefer to use the 2016 version, which is well tested and analysed in various studies. For details of DICE, we refer to Nordhaus (2017) and the references therein. This model maximises the social welfare function
@@
where $C_{t}$ is the consumption and $L_{t}$ is the population in the world in billions at time $t$, $U(C_{t},L_{t})=L_{t}(C_{t}/L_{t})^{1-\alpha}/(1-\alpha)$ is the utility per period, $\alpha=1.45$ is the risk aversion parameter, $\rho$ is a utility discounting factor; and time index $t=0,1,...$ corresponds to $\Delta=5$-year time step. This deterministic DICE-2016 model features six state variables: world produced economic capital $K_{t}$ in trillions of USD as of 2010, carbon concentrations (in billions of tons) in the atmosphere, the upper oceans, and the lower oceans $M_{t}^{\mathrm{AT}}$, $M_{t}^{\mathrm{UP}}$ and $M_{t}^{\mathrm{LO}}$ respectively, the global mean surface temperature $T_{t}^{\mathrm{AT}}$, and the mean temperature of the deep oceans $T_{t}^{\mathrm{LO}}$ (both temperatures are measured in degrees of C above the temperature during the year 1900).
@@
where $\delta_{K}$ is depreciation parameter and $Q_{t}=\Omega_{t}Y_{t}$ is the net economic annual output in a period $t$. The gross output is modelled by Cobb–Douglas production function of capital, labor, and technology
@@
where the second term is abatemant cost and the third term is the damage cost both as the fraction of the output; $\theta_{2}=2.6$, $\pi_{2}=0.00236$ and $\theta_{1,t}$ are parameters specified in DICE. The evolutions of carbon concentration $M_{t}=(M_{t}^{\mathrm{AT}},M_{t}^{\mathrm{UP}},M_{t}^{\mathrm{LO}})^{\top}$ and global temperature $T_{t}=(T_{t}^{\mathrm{AT}},T_{t}^{\mathrm{LO}})^{\top}$ are modelled as
@@
where $F_{t}=\eta\log_{2}({M_{t}^{\mathrm{AT}}}/{{M}_{\ast}^{\mathrm{AT}}})+F_{t}^{\mathrm{EX}}$ is radiative forcing, ${M}_{\ast}^{\mathrm{AT}}=588$ is pre-industrial carbon concentration, and $F_{t}^{\mathrm{EX}}$ is external forcing function. $\Phi_{M}$ is a 3 by 3 matrix of coefficients for the evolution of concentrations and $\Phi_{T}$ is a 2 by 2 matrix of coefficients for the evolution of temperatures. Here, we also note that $\beta=3.666$ represents the $\mathrm{CO}_{2}$ to carbon mass transformation coefficient, i.e., under the model parameterisation, emission is measured in tons of carbon dioxide while concentration is in tons of carbon. For a detailed discussion and all parameter values of the DICE model, we refer to Nordhaus (2018) and the GAMS code of DICE-2016. The functions $L_{t}$, $A_{t}$, $\sigma_{t}$, $E^{\mathrm{Land}}$ and $F_{t}^{\mathrm{EX}}$ are exogenous functions estimated in the original DICE2016 using various data sources, see Nordhaus (2018), different from SSP-IAM data.
The DICE model finds optimal carbon emission control $\mu_{t}$ and optimal consumption $C_{t}$ for each $t$ maximising the welfare function (1) with infinity time horizon replaced by $N=100$:
@@
where $\bm{C}=(C_{0},C_{1},...,C_{N})$ and $\bm{\mu}=(\mu_{0},\mu_{1},...,\mu_{N})$ are controls and $X_{t}$ is the vector of state variables $X_{t}=(K_{t},M_{t}^{\mathrm{AT}},M_{t}^{\mathrm{UP}},M_{t}^{\mathrm{LO}},T_{t}^{\mathrm{AT}},T_{t}^{\mathrm{LO}})^{\top}$ evolving in time according to equations (3,6,7) that can be written as a deterministic Markov decision process $X_{t+1}=\mathcal{T}_{t}(X_{t},C_{t},\mu_{t})$. This is a standard optimal control problem that can be solved using the Bellman equation backward in time for $t=N,N-1,...,0$ (see e.g. Bäuerle and Rieder, 2011)
@@
The orignal DICE solution in GAMS and Excel is brute force maximization in (8) with respect to $C_{0},...,C_{N}$ and $\mu_{0},...,\mu_{N}$.
@@
The second formula is used in the original GAMS code implementation of DICE with derivatives calculated numerically using small perturbation of the arguments in the welfare function from their optimal values; this representation is somewhat confusing in the original DICE literature, and we refer the reader to Braun et al. (2024) for precise mathematical formulation details. In our paper, we use R-code implementation of DICE producing identical results to the GAMS code subject to computational precision [^5] Our code is based on adapting the R-code https://github.com/olugovoy/climatedice.. For computational efficiency, we calculate $SCC_{t}$ directly as the net present value of damages by calculating the total discounted economic damages of the additional ton of $\mathrm{CO}_{2}$ released in year $t$ (the additional ton of $\mathrm{CO}_{2}$ yields reduced economic consumption in the future as the result of a climate damage function). This gives virtually the same results for SCC as obtained by the original GAMS code (p.284 Nordhaus, 2014); see also Braun et al. (2024) for the numerical calculation of SCC using different methods.
In this study, instead of optimal emission control, we consider deterministic scenarios for $\mu_{t}$ and maximize the welfare function with respect to the consumption only. Given scenarios for $\mu_{t}$ and optimal consumption values $C_{t}$, the trajectories of the economic and state variables are calculated using (3,6,7).
@@
$E_{t}=(1-\widetilde{\mu}_{t})(\sigma_{t}Y_{t}+E_{t}^{\mathrm{Land}}),$
i.e. $E_{t}=0$, when $\widetilde{\mu}_{t}=1$. The original emission control under this parameterisation is calculated as $\mu_{t}=\widetilde{\mu}_{t}\left(1+E_{t}^{\mathrm{Land}}/(\sigma_{t}Y_{t})\right)$.

### v2

# Scenario-based actuarial climate risk assessment via calibration of the DICE model to the shared socioeconomic pathways

@@ Abstract
Accounting for climate-related risks is an emerging problem for life insurers around the world. In this paper, we demonstrate how scenario trajectories for global temperature can be obtained using the cost-benefit Dynamic Integrated Climate-Economy (DICE) model calibrated to the five Shared Socioeconomic Pathways (SSPs). These scenarios can also be calculated under different carbon emission mitigation targets such as achieving net-zero carbon emissions by a specific year. We show how to calibrate the DICE model to align industrial and land-use carbon emissions with projections from six leading process-based integrated assessment models (IAMs): IMAGE, MESSAGE–GLOBIOM, AIM/CGE, GCAM, REMIND–MAgPIE and WITCH–GLOBIOM. The obtained scenario trajectories of global temperature can be linked to the climate-change induced excess mortality in various regions that, in turn, can be used for stress testing of life insurance portfolios. We illustrate this using synthetic portfolios of life insurance and annuity products.
Keywords: Actuarial climate risk, climate scenarios, life insurance, integrated climate-economy modelling, shared socioeconomics pathways, net-zero emission scenarios
@@ 1 Introduction
Global warming prompted the negotiation of the Paris Agreement in 2015 by 196 parties United Nations Treaty Collection (2015) aiming to limit the global temperature increase to $2^{\circ}\text{C}$ above preindustrial levels, with a preferred target of $1.5^{\circ}\text{C}$. To achieve these goals, many countries, cities, businesses, and other institutions have committed to achieving net-zero carbon emissions. Hereafter, “carbon emission” refers to the emission of all greenhouse gases. Over 140 countries, including China, India, and the European Union countries, have set goals to reach net-zero emissions by around 2050-2070. This has resulted in more public and private funds being provided for green projects involving renewable energy and energy conservation, as the world works to mitigate the effects of global climate change.
The impact of climate change on life and non-life insurance portfolios becomes an important area of risk management. Climate change increases not only the frequency and severity of weather-related non-life insurance claims but also mortality rates affecting life insurance products. Stress testing is a common practice of risk management undertaken by insurance companies, but is little studied with respect to climate change in the academic literature. Our paper demonstrates how stress test scenarios for temperature trajectories can be obtained and linked to temperature-related excess mortality for stress testing of life insurance portfolios.
A number of actuarial societies have developed the so-called climate indices [^1] See for example, The Actuaries Climate Indexes for USA and Canada, and Australia https://actuariesclimateindex.org/ and https://www.actuaries.asn.au/microsites/climate-index/ measuring the observed changes in extreme weather and sea levels to inform actuaries about climate trends and impacts related to non-life insurance. At the same time, climate change can significantly impact mortality rates by exacerbating heat-related, cardiovascular and respiratory illnesses. Of course, the impact of climate change on mortality depends on geographic location and death rate may even decrease in some cold climate countries. However, overall the increase in mortality on every continent is confirmed by several studies, see e.g. Vicedo-Cabrera et al. (2021); Gasparrini et al. (2017).
The European Insurance and Occupational Pensions Authority (EIOPA) put forward stress testing principles for insurers EIOPA (2022) that provides high level guidance to insurers. The most recent application paper by the International Association of Insurance Supervisors IAIS (2025) discusses the key topics relevant for the supervision of climate-related risks including stress test/scenario analysis. It defines climate-related risk (page 10) as “risk posed by the exposure of an insurer to physical, transition and/or litigation risks caused by or related to climate change.” It also defines (page 75) scenario analysis as “A method of assessment that considers the impact of a combination of circumstances to reflect historical or other scenarios which are analysed in light of current conditions. Such analysis may be conducted deterministically or stochastically”. It recommends to consider a very wide range scenarios for global emissions including net-zero scenarios. Although it does not prescribe specific scenarios, it mentions scenarios developed by the Network of Central Banks and Supervisors for Greening the Financial System (NGFS) [^2] https://www.ngfs.net/, see NGFS (2024). These scenarios are based on several process-based Integrated Assessment Models (IAMs) and Shared Socioeconomic Pathways (SSPs).
@@
Since the early 1990s, IAMs have become a central analytical tool in climate change economics; see reviews in Weyant (2017). The mitigation and climate impact assessments in the Intergovernmental Panel on Climate Change (IPCC) Assessment Reports [^3] The IPCC is the United Nations body responsible for assessing the science related to climate change. Its latest report, the Sixth Assessment Report (2023), is available at www.ipcc.ch. have primarily relied on IAMs developed by various research groups. These models can be broadly categorised into two types: cost-benefit IAMs, which estimate optimal mitigation strategies by balancing economic costs and climate damages, and process-based IAMs, which form the basis of IPCC’s assessments of transformation pathways towards temperature targets. However, cost-benefit IAMs have historically relied on a single reference optimal pathway and have not systematically incorporated the full range of Shared Socioeconomic Pathway (SSP) scenarios. As a result, there is a lack of integrated research pursued in our paper that quantitatively aligns the baseline assumptions of process-based IAMs with cost-benefit modelling frameworks under alternative net-zero emission trajectories.
@@
In this paper, we develop scenarios of global temperature and use these scenarios to find projections of mortality rates driven by changes in the temperature that affect life insurance products. We study this using the cost-benefit dynamic integrated climate-economy (DICE) model calibrated to baseline SSPs and emissions from the leading process-based models. DICE was originally proposed more than 30 years ago in Nordhaus et al. (1992) and was regularly refined with the latest beta version update in 2023444The most recent version of the DICE model is available at https://williamnordhaus.com.. It is a cost-benefit IAM that estimates optimal mitigation levels relative to the economic costs of climate impacts. DICE model is one of the three main IAMs used by the United States government to determine the social cost of carbon; see Interagency Working Group on Social Cost of Greenhouse Gases (2016). The other two are PAGE (Policy Analysis of the Greenhouse Effect model, see Hope (2008)) and FUND (Climate Framework for Uncertainty, Negotiation, and Distribution model, see Anthoff and Tol (2010)). Even though the DICE model has limitations in the model structure and model parameters debated in the literature Grubb et al. (2021); Pindyck (2017), it has become the typical reference point for climate-economy modelling, and is used in our study. The DICE model is a global model but fits for purpose with respect to modelling global temperature and global carbon concentration, though it may need recalibration of model parameters in view of recent findings in Folini et al. (2025). We also note that, regional version of DICE (RICE model) is available Nordhaus and Yang (1996) that allows for multiple regions to model region specific economies but its climate model is the same as in DICE.
@@
Almost any process-based IAM can be run for different SSPs. Among these, for each SSP a single IAM interpretation was selected as the so-called representative marker scenario for recommended use by future analyses of climate change, its impacts, and response measures. These are referred to as marker SSPs, discussed in detail in e.g. Riahi et al. (2017) and defined in Section 2.3. For example, marker SSP1 corresponds to SSP1 run using IMAGE. The selection of markers was guided by two main considerations: the internal consistency of the full set of SSP markers, and the ability of the different models to represent distinct characteristics of the storylines. The SSP data and process-based IAMs forecasts are available from SSPdata-IIASA. This database does not include cost-benefit IAMs such as DICE. Even though marker SSPs are standard references, non-marker SSPs are also considered to be important since they provide insights into possible alternative scenario interpretations of the same basic SSP elements and storylines. In our study, we calculate results for DICE model calibrated to baseline scenarios of all five marker SSPs; and we also calculate non-marker scenarios when IAM data are available for all five SSPs.
@@
In our study, we re-calibrate the DICE model to match the SSP baseline marker and non-marker scenarios [^6] The baseline SSP scenarios are the reference cases for mitigation, climate impacts and adaptation analyses. They represent future developments in the absence of new climate policies and mitigation scenarios. under the main six process-based IAMs for the emission of greenhouse gases. Then, re-calibrated DICE model is used to produce scenarios of future for climate and economy variables under SSP1–SSP5 and assuming scenarios for carbon emission reduction leading to net-zero carbon emission in 2050 and 2100 (we also calculated scenarios for zero industrial emission only). One of the main observations from these projections is that even if a zero emission target is achieved worldwide by 2050, the temperature increased by 2100 will be larger than 2C in the range 2.5-2.7C. For this scenario, the social cost of carbon is increasing about 10 times from USD 30-50 in 2025 to USD 250-400 in 2100 depending on SSP-IAM. This provides a carbon price benchmark for policy makers and is generally consistent with Riahi et al. (2022) and Yang et al. (2018).
If temperature-related excess mortality is estimated, then temperature projections translate into scenarios for mortality rates, that in turn impact insurance premia and reserves. In this paper, we use climate-induced mortality from Bressler (2021). Note that for illustration we use global temperature related excess mortality function, but for practical use, one should use country specific excess mortality functions such as estimated in Lee et al. (2020) or Gasparrini et al. (2017). The scenarios based on the DICE model developed in our paper should be considered as complementing NGFS scenarios developed using three process-based models (REMIND-MAgPIE, MESSAGE-GLOBIOM, and GCAM). The cost-benefit DICE model maximises the welfare function wrt consumption and emission controls, i.e. allows for adaptive optimal decisions rather than fixed scenarios under the process-based models.
@@ 2 Methodology
In this section, we outline the main DICE model equations, the calibration of the DICE exogenous functions to the SSP-IAM data, and the excess temperature-related mortality function. The results for climate-economy trajectories and the impoacts on life insurance products under DICE calibrated to different SSPs-IAMs will be presented in the next section.
@@ 2.1 Model
In our study, we calculate the climate-economy projections using DICE-2016. Specifically, we use the version DICE2016R-091916ap.gms. We note that beta 2023 version is available, but we prefer to use the 2016 version, which is well tested and analysed in various studies. For details of DICE, we refer to Nordhaus (2017) and the references therein. This model maximises the social welfare function
@@
where $C_{t}$ is the consumption and $L_{t}$ is the population of the world in billions at time $t$, $U(C_{t},L_{t})=L_{t}(C_{t}/L_{t})^{1-\alpha}/(1-\alpha)$ is the utility per period, $\alpha=1.45$ is the risk aversion parameter, $\rho$ is a utility discounting factor; and time index $t=0,1,$... corresponds to $\Delta=5$-year time step. This deterministic DICE-2016 model features six state variables: world produced economic capital $K_{t}$ in trillions of USD as of 2010, carbon concentrations (in billions of tons) in the atmosphere, the upper oceans, and the lower oceans $M_{t}^{\mathrm{AT}}$, $M_{t}^{\mathrm{UP}}$ and $M_{t}^{\mathrm{LO}}$ respectively, the global mean surface temperature $T_{t}^{\mathrm{AT}}$, and the mean temperature of the deep oceans $T_{t}^{\mathrm{LO}}$ (both temperatures are measured in degrees of C above the temperature during the year 1900).
@@
where $\delta_{K}$ is the depreciation parameter and $Q_{t}=\Omega_{t}Y_{t}$ is the net economic annual output in a period $t$. The gross output $Y_{t}$ is modelled by the Cobb–Douglas production function of capital, labor, and technology
@@
where the second term is the abatement cost and the third term is the damage cost both as the fraction of the output; $\theta_{2}=2.6$, $\pi_{2}=0.00236$ and $\theta_{1,t}$ are parameters specified in DICE. The evolutions of carbon concentration $M_{t}=(M_{t}^{\mathrm{AT}},M_{t}^{\mathrm{UP}},M_{t}^{\mathrm{LO}})^{\top}$ and global temperature $T_{t}=(T_{t}^{\mathrm{AT}},T_{t}^{\mathrm{LO}})^{\top}$ are modelled as
@@
where $F_{t}=\eta\log_{2}({M_{t}^{\mathrm{AT}}}/{{M}_{\ast}^{\mathrm{AT}}})+F_{t}^{\mathrm{EX}}$ is the radiative forcing, ${M}_{\ast}^{\mathrm{AT}}=588$ is the pre-industrial carbon concentration, and $F_{t}^{\mathrm{EX}}$ is the external forcing function. $\Phi_{M}$ is a 3 by 3 matrix of coefficients for the evolution of concentrations and $\Phi_{T}$ is a 2 by 2 matrix of coefficients for the evolution of temperatures. Here, we also note that $\beta=3.666$ represents the $\mathrm{CO}_{2}$ to carbon mass transformation coefficient, i.e., under the model parameterisation, emission is measured in tons of carbon dioxide while concentration is in tons of carbon. For a detailed discussion and all parameter values of the DICE model, we refer to Nordhaus (2018) and the GAMS code of DICE-2016. In the original DICE-2016, functions $L_{t}$, $A_{t}$, $\sigma_{t}$, $E^{\mathrm{Land}}$ and $F_{t}^{\mathrm{EX}}$ are exogenous functions estimated using various data sources (see Nordhaus (2018)), different from the SSP-IAM data.
In this paper we follow Bressler (2021) extending the DICE model treating population as an additional state variable that also depends on temperature:
$L_{t+1}$ $=L_{t}\times\left(1+b_{t}-d_{t}(1+\delta(T_{t}^{\mathrm{AT}}))\right)$
$=L_{t}\times(1+b_{t}-d_{t})-N_{t}\times\delta(T^{\mathrm{AT}}_{t}),$ (8)
where $b_{t}$ is the fertility rate and $d_{t}$ is the mortality rate. For numerical illustration, we will use the global excess mortality function $\delta(x)=ax^{\nu}$ with $a=0.0001811$ and $\nu=3.745$ estimated in Bressler (2021). The factor $(1-b_{t}-d_{t})$ can be calibrated using SSP data as $(1-b_{t}-d_{t})=\hat{L}_{t+1}/{\hat{L}_{t}}$, where $\hat{L}_{t}$ is population under the corresponding SSP, and the number of deaths $N_{t}=L_{t}d_{t}$ is estimated as the number of deaths projected under different SSPs in (Sellers, 2020, Figure 2).
The DICE model finds optimal carbon emission control $\mu_{t}$ and optimal consumption $C_{t}$ for each $t$ maximising the welfare function (1) with an infinite time horizon replaced by $N=100$:
@@
where $\bm{C}=(C_{0},C_{1},...,C_{N})$ and $\bm{\mu}=(\mu_{0},\mu_{1},...,\mu_{N})$ are controls and $X_{t}$ is the vector of state variables $X_{t}=(K_{t},M_{t}^{\mathrm{AT}},M_{t}^{\mathrm{UP}},M_{t}^{\mathrm{LO}},T_{t}^{\mathrm{AT}},T_{t}^{\mathrm{LO}},L_{t})^{\top}$ evolving in time according to equations (3,6,7) that can be written as a deterministic Markov decision process $X_{t+1}=\mathcal{T}_{t}(X_{t},C_{t},\mu_{t})$. This is a standard optimal control problem that can be solved using the Bellman equation backward in time for t=N,N1,,0t=N,N-1,...,0 (see e.g. Bäuerle and Rieder, 2011)
@@
The orignal DICE solution in GAMS and Excel is brute force maximisation in (9) with respect to $C_{0},...,C_{N}$ and $\mu_{0},...,\mu_{N}$.
@@
The second formula is used in the original GAMS code implementation of DICE with derivatives calculated numerically using small perturbation of the arguments in the welfare function from their optimal values; this second representation is somewhat confusing in the original DICE literature, and we refer the reader to Braun et al. (2024) for precise mathematical formulation details. In our paper, we use the R-code implementation of DICE producing identical results to the GAMS code subject to computational precision [^7] Our code is based on adapting the R-code https://github.com/olugovoy/climatedice.. For computational efficiency, we calculate $SCC_{t}$ directly as the net present value of damages by calculating the total discounted economic damages of the additional ton of $\mathrm{CO}_{2}$ released in year $t$ (the additional ton of $\mathrm{CO}_{2}$ yields reduced economic consumption in the future as the result of a climate damage function). This gives virtually the same results for SCC as obtained by the original GAMS code (p.284 Nordhaus, 2014); see also Braun et al. (2024) for the numerical calculation of SCC using different methods.
In our paper, we also consider deterministic scenarios for $\mu_{t}$, that requires maximisation of the welfare function with respect to consumption only. Given scenarios for $\mu_{t}$ and optimal consumption values of $C_{t}$, the trajectories of the economic and state variables are calculated using (3,6,7).
@@
$E_{t}=(1-\widetilde{\mu}_{t})\times\left(\sigma_{t}Y_{t}+E_{t}^{\mathrm{Land}}\right),$
i.e. $E_{t}=0$, when $\widetilde{\mu}_{t}=1$. The original emission control under this parameterisation is calculated as $\mu_{t}=\widetilde{\mu}_{t}\times\left(1+E_{t}^{\mathrm{Land}}/(\sigma_{t}Y_{t})\right)$.
2.2 Climate change induced excess mortality
Mortality rate projections is the main input for life insurance products and reserves. Various deterministic and stochastic mortality models have been studied in the literature Booth and Tickle (2008). Denote the baseline annual probability of death at age $x$ in year $\tau$ by $q_{x}(\tau)$, then the mortality rate accounting for the temperature-related excess mortality can be written as
$q^{\mathrm{adj}}_{x}(\tau)=q_{x}(\tau)\times\big(1+\widetilde{\delta}(T^{\mathrm{AT}}(\tau))\big).$
[…]
${{}_{k}}p_{x}^{\mathrm{adj}}(\tau)=\prod_{j=0}^{k-1}(1-q_{x+j}^{\mathrm{adj}}(\tau+j)).$ (11)
Once scenario realisation of $T_{t}^{\mathrm{AT}}$ is known for $t=0,1,$..., then $\widetilde{\delta}(T^{\mathrm{AT}}(\tau))$ can easily be evaluated for each future year $\tau$. Practitioners can use available temperature projections from various process based IAMs or/and NGFS scenarios. In our paper, we develop scenarios based on the cost-benefit DICE model that maximises the welfare function wrt consumption and emission controls, i.e. allows for adaptive optimal decisions rather than fixed scenarios under the process-based models.
[…]